OpenStudy (anonymous):

Your mortgage is $135,400.00 at 5.25% APR. If you pay $1,500.00 toward the mortgage each month, how much of your second month's payment is applied to the principal?

OpenStudy (phi):

I would find the interest that is due after one month. change 5.25 % APR (i.e. per year) to 0.0525/12 (interest rate per month) multiply that number times the mortgage 135400 and add this to the original amount... or... multiply 135400 by (1+0.0525/12) to get the new amount. now subtract off the payment 1500. what is left is the new amount you owe. multiply that amount by 0.0525/12 to get the interest that you add. the payment 1500 will pay off that interest and what is left over will pay some of the principal . that amount left over is the answer to the question.

OpenStudy (anonymous):

@phi I think I keep messing up at some point, I can't seem to get an answer that's on the list of answers I can choose from.

OpenStudy (phi):

what do you get for the interest for the first month ?

OpenStudy (anonymous):

The interest? 592.375, which I rounded up to 592.38.

OpenStudy (phi):

ok. when we make the first payment of 1500 that covers the interest, and leaves 907.63 left to pay the principal that means after the first payment, the principal is $135,400.00 - 907.63 = 134,492.38 ok so far ?

OpenStudy (anonymous):

Yeah, I think I understand that.

OpenStudy (phi):

or we could just say: at the end of 1 month we owe 135400+592.38= 135,992.38 then make the payment of 1500 and now owe 135,992.38-1500 = 134,492.38

OpenStudy (anonymous):

Alright, yeah I understand that.

OpenStudy (phi):

next, what is one month's interest on 134,492.38

OpenStudy (anonymous):

588.40?

OpenStudy (phi):

yes. now we make the payment. we pay that interest first 1500-588.40 = 911.5958 or 911.60 left over to pay down the principal hopefully that is one of your choices.

OpenStudy (anonymous):

Yeah, it is! I see where I messed up now, as well. Thank you very much. I really appreciate your help.